Find an equation of the line containing the point (5,2) perpendicular to the line 3x-5y=10.
Answer by Cromlix(4381)

Sort 3x - 5y = 10 into y = mx + c form.
-5y = -3x + 10
5y = 3x - 10
y = 3/5x - 2
Lines that are perpendicular to one
another have gradients that multiply
together to give -1
m1 * m2 = -1
The above line has a gradient of 3/5
so, line perpendicular to it has a
gradient of -5/3
Using point (5,2) and m = -5/3
y - b = m(x - a)
y - 2 = -5/3(x - 5)
y - 2 = -5/3x + 25/3
y = -5/3x + 25/3 + 6/3
y = -5/3x + 31/3
OR multiply thro' by 3
3y = -5x + 31.
Hope this helps.
:-)